OpenStudy (anonymous):

For the function f(x) = (8-2x)2 ,find f-1 . Determine whether f-1 is a function. A. f-1(x) = ;f-1 is a function B. f -1 (x) = ; f-1 is not a function C. f-1(x) = ;f-1 is not a function D. f-1(x) = ; f-1 is a function Its not A!

OpenStudy (anonymous):

@KryoWolf @DJ♬Øϟϟ¥ @Shay17

OpenStudy (anonymous):

im here replace each (x) with the number 1

OpenStudy (anonymous):

whats the difference between D and A

OpenStudy (anonymous):

A function is 1 to 1; meaning, one input will result in exactly 1 output. A counterexample would be where the input results in two possible outputs. An example of this would be \(f(x) = (\sqrt{x})\). If we input 4, the answers are plus or minus 2, since +2 squared = 4 and -2 squared = 4.

OpenStudy (anonymous):

I don't think there is one that's how some of them have been coming out

OpenStudy (anonymous):

ok well, just replace all (x)s with 1,2,3, ect.

OpenStudy (anonymous):

okay then since there are really only 2 choices how would that work?

OpenStudy (anonymous):

You need to find the inverse first. To do this, substitute f(x) with y, and solve for x. Once you solve for x, replace x with \(f^{-1}(x)\) and replace y with x.

OpenStudy (anonymous):

so it B?

OpenStudy (anonymous):

did you do the math to figure it out?

OpenStudy (anonymous):

if so, show me your work

OpenStudy (anonymous):

no I kinda just put it all together did what yall said and since a&c are the exact and so is b&d I chose b:/ it kinda confusing me

OpenStudy (anonymous):

\[f(x) = (8-2x)^2 \rightarrow y = (8-2x)^2 \rightarrow \sqrt{y} = 8-2x \]\[\rightarrow \sqrt{y} -8 = 2x\rightarrow\frac{\sqrt{y}-8}{2} = x \rightarrow f^{-1}(x) = \frac{\sqrt{x}-8}{2}\]

OpenStudy (anonymous):

@Shay17 #16

OpenStudy (anonymous):

Now, you see that since you input values for x into your inverse function, you get + or - x values because of the square-root, as I explained earlier. So, this is not a function.

OpenStudy (anonymous):

well, im not sure about this question either, so i'll let digitallogic take over

OpenStudy (anonymous):

okay so it is B? :)

OpenStudy (anonymous):

I don't know, because you didn't show your answer choices completely. Pick the one that looks like the inverse function I gave you.

OpenStudy (anonymous):

as far as i know, yes

OpenStudy (anonymous):

what you see is what I see that what it looks like

OpenStudy (anonymous):

Your answers don't contain the inverse functions?

OpenStudy (anonymous):

@digitallogic can you help with the question i have?

OpenStudy (anonymous):

Take a screenshot. This is what I see that you posted. The functions are missing.

OpenStudy (anonymous):

nope just a blank space

OpenStudy (anonymous):

That's stupid. Your test is messed up. The answer is B or C but you can't tell because there are no function values. Complain to your teacher for being dumb.

OpenStudy (anonymous):

Because B and C are identical choices without the functions.

OpenStudy (anonymous):

digitallogic can u help with my question?

OpenStudy (anonymous):

Depends on the subject @KryoWolf

OpenStudy (anonymous):

lol will do :) but thanks for the help and right that why I was confused

OpenStudy (anonymous):

Cool :) No problemo.

OpenStudy (anonymous):

@digitallogic its linear models